# RGB-to-Polarization Estimation: A New Task and Benchmark Study

### v1

@@ 1 Introduction
Polarization images contain rich physical information that is not captured by standard RGB cameras, such as birefringence, surface stress, roughness, and other material properties wolff1990polarization . This polarization information provides valuable visual cues that enhance a variety of computer vision tasks, such as reflection separation nayar1997separation ; lei2020polarized ; pang2022progressive , material analysis tasks like classification and segmentation wolff1990polarization ; mei2022glass ; liang2022multimodal , and shadow removal zhou2025polarization .
@@
In response to this question, we introduce a new task: RGB-to-polarization image estimation. As illustrated in Figure 1, the goal is to leverage neural networks to estimate polarization information directly from RGB inputs. To the best of our knowledge, this is the first work that explores sensor-free polarimetric imaging using neural networks. In this work, polarization information is represented using the Stokes parameters. We treat the RGB image—corresponding to the total intensity ($\mathrm{S}_{0}$)—as input, and train models to estimate the remaining polarization components, $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$, which respectively characterize horizontal/vertical linear polarization ($0^{\circ}$/$90^{\circ}$), diagonal linear polarization (±$\pm 45^{\circ}$), and circular polarization. These components are essential for a wide range of downstream polarization-based vision tasks.
Building on this task, we establish a comprehensive benchmark for RGB-to-polarization image estimation. Leveraging one of the latest RGB-polarization datasets jeon2024spectral , we standardize evaluation protocols to ensure consistent and comparable results across methods. Multiple deep learning models, spanning both restoration-based and generative architectures, are further evaluated to assess their effectiveness on this task. Through systematic analysis, our benchmark reveals the strengths and limitations of existing approaches and offers insights for future development in sensor-free polarization estimation. Our contributions are summarized as follows:
@@ 2 Related work
Existing image restoration backbones liang2021swinir ; wang2022uformer ; zamir2022restormer ; chen2022simple ; li2023efficient ; zamir2021multi are widely used in low-level vision tasks and can be naturally adapted for RGB-to-polarization image estimation. Transformer-based architectures such as SwinIR liang2021swinir , Uformer wang2022uformer , and Restormer zamir2022restormer leverage self-attention mechanisms to capture long-range dependencies, achieving strong performance across denoising, deraining, and deblurring tasks. In our benchmark, we adopt Uformer and Restormer as representative restoration backbones due to their effectiveness and generalizability.
@@
Existing generative models, such as GANs goodfellow2014generative and diffusion models ho2020ddpm , have demonstrated strong capabilities in modeling complex image distributions. In this work, we explore whether diffusion models can be used to estimate polarization images from RGB inputs. Unlike restoration-based models such as Restormer, which perform direct pixel-wise regression, diffusion models learn to approximate the underlying data distribution and generate outputs through iterative denoising. This generative formulation allows them to capture uncertainty and produce more flexible predictions. Following our task setting, we adopt two representative diffusion models, WDiff ozdenizci2023restoring and DiT peebles2023scalable , both trained in a conditional setting where the RGB image guides the generation of polarization components.
@@
When extended to computer vision tasks, polarization plays a critical role in distinguishing surface properties. Specifically, the polarization state of reflected light changes depending on factors such as a material’s refractive index, surface roughness, coating, or texture—enabling the classification of materials and object recognition wolff1990polarization . In addition, specular reflections can be suppressed by exploiting polarization properties, and the contrast of transparent or translucent objects can be enhanced using a polarizer riviere2017polarization . Furthermore, materials exhibiting birefringence or optical activity cause characteristic changes in polarization, often reflected in the $\mathrm{S}_{2}$ and $\mathrm{S}_{3}$ components, which are useful for biomedical imaging, tissue analysis, and fiber inspection deboer2002review . Finally, surface normals affect the polarization state of reflected light, which in turn provides geometric cues for shape or depth inference in tasks such as Shape-from-Polarization (SfP) based 3D reconstruction kadambi2015polarized . These cues are absent in standard RGB images, highlighting the unique and valuable role of polarization information.
@@
Figure 2: (a) The Poincaré sphere serves as a geometric representation for describing all possible states of polarization using the Stokes parameters. (b–d) Polarization describes how the electric field of a light wave oscillates within the plane perpendicular to the direction of propagation. (b) Linear polarizations at 0° and 90°. (c) Linear polarizations at ±$\pm 45^{\circ}$ angles. (d) Right- and left-handed circular polarizations.
@@ 3 Method
3.1 Stokes parameters
The polarization state of light characterizes how the electric field oscillates within the plane perpendicular to the direction of propagation. In this work, we represent the polarization state of light using the Stokes vector and visualize it on the Poincaré sphere, as illustrated in Figure 2, where each point corresponds to a unique polarization state. A light wave’s polarization can be described using the four Stokes parameters $(\mathrm{S}_{0},\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3})$, where $\mathrm{S}_{0}$ denotes the total intensity of the beam, and the remaining components collectively characterize the polarization state. Specifically, $\mathrm{S}_{1}$ represents the difference in intensity between 0 and 90 linear polarization; $\mathrm{S}_{2}$ corresponds to the difference between +45 and -45 linear polarization; and $\mathrm{S}_{3}$ indicates the difference between right- and left-handed circular polarization mcmaster1954polarization ; schaefer2007measuring .
@@ 3.1 Stokes parameters
where $\mathrm{S}_{0}$ denotes the total light intensity, represented as an RGB image. On the Poincaré sphere, the polarization direction is represented by the vector $\mathbf{S}=(\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3})/\mathrm{S}_{0}$. For instance, $\mathbf{S}=(1,0,0)$ corresponds to 0 linear polarization, $(-1,0,0)$ to 90 linear polarization, $(0,1,0)$ and $(0,-1,0)$ to ±45 linear polarization, and $(0,0,\pm 1)$ to right- or left-handed circular polarization. Elliptical polarization states lie between these extremes, depending on the value of $\mathrm{S}_{3}$. Each component of the Stokes vector carries distinct material-sensitive information: high $\mathrm{S}_{1}$ values often indicate smooth dielectric or metallic surfaces due to dominant specular reflection; elevated $\mathrm{S}_{2}$ suggests birefringent or fibrous materials; and strong $\mathrm{S}_{3}$ values arise in scattering or optically active media, such as biological tissues or rough dielectrics riviere2017polarization .
@@ 3.2 Benchmark Design
Next, we build our benchmark on the recent large-scale RGB-polarization dataset proposed by Jeon et al. jeon2024spectral . It provides high-quality, spatially aligned image pairs consisting of RGB inputs $\mathrm{S}_{0}$ and their corresponding Stokes components $[\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3}]$, enabling supervised training. All images are resized to a fixed resolution of $H\times W$ before training. Both the RGB inputs and the Stokes targets are normalized to the range $[0,1]$ for consistent training across models.
@@
Following standard practices in low-level vision, we evaluate model performance using three complementary metrics: peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS). PSNR and SSIM assess pixel-level accuracy and structural fidelity, while LPIPS captures perceptual similarity in the feature space. All metrics are computed independently for each Stokes component and then averaged to report the overall performance.
@@ 4 Experiments
In our experiments, we use the dataset provided by Jeon et al. jeon2024spectral for training and evaluation. Specifically, the first 1,000 RGB–Stokes image pairs are used for training, and the last 200 pairs from the dataset are reserved for testing. For a fair comparison, all image pairs are resized to 256$\times$256 for training and testing across all baseline models.
@@ 4.1 Implementation details
All experiments are conducted on a server equipped with four NVIDIA RTX A5000 GPUs, each with 24 GB of memory. To ensure reproducibility, we will release all code, model checkpoints, and configuration files upon acceptance. The following parts detail the network architectures and training settings used across all evaluated models.
@@
The training steps for the three networks are set to 70,000, 70,000, and 40,000, respectively. Due to differences in patch size and network depth, we adopt different batch sizes for training. For each model, we use the maximum batch size that fits within GPU memory constraints. The batch sizes are set to 16, 48, and 128, respectively. During training, we use the Adam optimizer with an initial learning rate of 1.5$\times$10-4.
@@
Generation baselines
The training steps for WDiff, DiT, RealFill, and Img2ImgTurbo are set to 90,000, 90,000, 10,000, and 8,000, respectively. We use the Adam optimizer with an initial learning rate of 1.5$\times$10-4 for WDiff and DiT, and adopt batch sizes of 128 and 258, respectively. For RealFill and Img2ImgTurbo, due to higher memory consumption, the batch sizes are limited to 12 and 4, respectively. To simulate a larger effective batch size, we apply gradient accumulation with a factor of 4 for Img2ImgTurbo. Following tang2024realfill , we adopt separate learning rates for LoRA modules in RealFill: 2e-4 for the UNet and 4e-5 for the text encoder. For Img2ImgTurbo parmar2024one , we set a unified learning rate of 5e-6 for all learnable parameters.
@@ 4.3 Qualitative evaluation
Figures 3, 4, 5, and 6 show qualitative results of the estimated Stokes components generated by DiT peebles2023scalable , Img2ImgTurbo parmar2024one , Uformer wang2022uformer , and MAE he2022masked . Each column visualizes the reconstructed $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$ components produced by each method. The results reveal diverse reconstruction characteristics across models, including differences in texture sharpness, structural consistency, and polarization pattern styles. Overall, MAE and Uformer demonstrate superior visual quality compared to diffusion-based approaches. MAE is particularly effective in preserving fine structural details and material boundaries, with minimal artifacts across all Stokes components. However, the restored details from restoration-based methods still deviate from the correct results, indicating that such models may still lack the capacity to capture the underlying polarization cues precisely. In contrast, diffusion-based methods struggle to reconstruct the correct appearance and structure of polarization components, indicating difficulty in modeling these signals. Although Img2ImgTurbo achieves higher PSNR scores than DiT, its visual results are less faithful to the ground truth, particularly in terms of structural detail. This suggests a trade-off between pixel-level accuracy and perceptual quality among diffusion-based models.

### v2

@@ 1 Introduction
Polarization images contain rich physical information that is not captured by standard RGB cameras, such as birefringence, surface stress, roughness, and other material properties wolff1990polarization . This polarization information provides valuable visual cues that enhance a variety of computer vision tasks, such as reflection separation nayar1997separation ; lei2020polarized ; pang2022progressive , material analysis tasks like classification and segmentation wolff1990polarization ; mei2022glass ; liang2022multimodal ; zhang2025mamba ; yang2025erf ; zhang2024heap , and shadow removal zhou2025polarization .
@@
In response to this question, we introduce a new task: RGB-to-polarization image estimation. As illustrated in Figure 1, the goal is to leverage neural networks to estimate polarization information directly from RGB inputs. To the best of our knowledge, this is the first work that explores sensor-free polarimetric imaging using neural networks. In this work, polarization information is represented using the Stokes parameters. We treat the RGB image—corresponding to the total intensity ($\mathrm{S}_{0}$)—as input, and train models to estimate the remaining polarization components, $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$, which respectively characterize horizontal/vertical linear polarization ($0^{\circ}$$90^{\circ}$), diagonal linear polarization ($\pm 45^{\circ}$), and circular polarization. These components are essential for a wide range of downstream polarization-based vision tasks.
Building on this task, we establish a comprehensive benchmark for RGB-to-polarization image estimation. Leveraging three of the latest RGB-polarization datasets jeon2024spectral ; qiu2021demos ; kurita2023sparse , we standardize evaluation protocols to ensure consistent and comparable results across methods. Multiple deep learning models, spanning both restoration-based and generative architectures, are further evaluated to assess their effectiveness on this task. Through systematic analysis, our benchmark reveals the strengths and limitations of existing approaches and offers insights for future development in sensor-free polarization estimation. It should be noted that multiple polarization states can correspond to the same RGB appearance, since RGB encodes only intensity and color but not the vectorial nature of light, though this ambiguity affects all methods equally and thus does not compromise the fairness of the benchmark. Our contributions are summarized as follows:
@@ 2 Related work
Existing image restoration backbones liang2021swinir ; wang2022uformer ; zamir2022restormer ; chen2022simple ; li2023efficient ; zamir2021multi ; chen2024dual ; lin2024nightrain ; lin2025nighthaze ; jin2023enhancing are widely used in low-level vision tasks and can be naturally adapted for RGB-to-polarization image estimation. Transformer-based architectures such as SwinIR liang2021swinir , Uformer wang2022uformer , and Restormer zamir2022restormer leverage self-attention mechanisms to capture long-range dependencies, achieving strong performance across denoising, deraining, and deblurring tasks. In our benchmark, we adopt Uformer and Restormer as representative restoration backbones due to their effectiveness and generalizability.
@@
Existing generative models, such as GANs goodfellow2014generative and diffusion models ho2020ddpm , have demonstrated strong capabilities in modeling complex image distributions. In this work, we explore whether diffusion models can be used to estimate polarization images from RGB inputs. Specifically, we adopt two representative conditional diffusion models, WDiff ozdenizci2023restoring and DiT peebles2023scalable , where the RGB image guides the generation of polarization components.
@@
When extended to computer vision tasks, polarization plays a critical role in distinguishing surface properties. Specifically, the polarization state of reflected light changes depending on factors such as a material’s refractive index, surface roughness, coating, or texture—enabling the classification of materials and object recognition wolff1990polarization . In addition, specular reflections can be suppressed by exploiting polarization properties, and the contrast of transparent or translucent objects can be enhanced using a polarizer riviere2017polarization . Furthermore, materials exhibiting birefringence or optical activity cause characteristic changes in polarization, often reflected in the $\mathrm{S}_{2}$ and $\mathrm{S}_{3}$ components, which are useful for biomedical imaging, tissue analysis, and fiber inspection deboer2002review . Finally, surface normals affect the polarization state of reflected light, which in turn provides geometric cues for shape or depth inference in tasks such as Shape-from-Polarization (SfP) based 3D reconstruction kadambi2015polarized . Beyond SfP, polarization has also been exploited for reflectance and appearance modeling baek2018polarimetric ; hwang2022sparse ; ha2024polarimetric , as well as for task-specific applications including polarization-aware semantic segmentation liu2025sharecmp , sparse polarization sensing kurita2023sparse , and wearable robotics yang2018polarization . These cues are absent in standard RGB images, highlighting the unique and valuable role of polarization information.
@@
Figure 2: (a) The Poincaré sphere serves as a geometric representation for describing all possible states of polarization using the Stokes parameters. (b–d) Polarization describes how the electric field of a light wave oscillates within the plane perpendicular to the direction of propagation. (b) Linear polarizations at 0° and 90°. (c) Linear polarizations at $\pm 45^{\circ}$ angles. (d) Right- and left-handed circular polarizations.
@@ 3 Method
3.1 Stokes Parameters and Derived Quantities
Stokes parameters.
The polarization state of light characterizes how the electric field oscillates within the plane perpendicular to the direction of propagation. In this work, we represent the polarization state of light using the Stokes vector and visualize it on the Poincaré sphere, as illustrated in Figure 2, where each point corresponds to a unique polarization statemcmaster1954polarization ; Yuan2025JLT ; schaefer2007measuring ; Yuan2025APL ; sheppard2014jones . A light wave’s polarization can be described using the four Stokes parameters $(\mathrm{S}_{0},\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3})$, where $\mathrm{S}_{0}$ denotes the total intensity of the beam, and the remaining components collectively characterize the polarization state. Specifically, $\mathrm{S}_{1}$ represents the difference in intensity between 0 and 90 linear polarization; $\mathrm{S}_{2}$ corresponds to the difference between +45 and -45 linear polarization; and $\mathrm{S}_{3}$ indicates the difference between right- and left-handed circular polarization mcmaster1954polarization ; schaefer2007measuring ; Yuan2025cleo .
@@ 3.1 Stokes Parameters and Derived Quantities
where $\mathrm{S}_{0}$ denotes the total light intensity, represented as an RGB image. On the Poincaré sphere, the polarization direction is represented by the vector $\mathbf{S}=(\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3})/\mathrm{S}_{0}$. For instance, $\mathbf{S}=(1,0,0)$ corresponds to 0 linear polarization, (1,0,0)(-1,0,0) to 90 linear polarization, $(0,1,0)$ and (0,1,0)(0,-1,0) to ±45 linear polarization, and $(0,0,\pm 1)$ to right- or left-handed circular polarization. Elliptical polarization states lie between these extremes, depending on the value of $\mathrm{S}_{3}$. Each component of the Stokes vector carries distinct material-sensitive information: high $\mathrm{S}_{1}$ values often indicate smooth dielectric or metallic surfaces due to dominant specular reflection; elevated $\mathrm{S}_{2}$ suggests birefringent or fibrous materials; and strong $\mathrm{S}_{3}$ values arise in scattering or optically active media, such as biological tissues or rough dielectrics riviere2017polarization .
Interpretable Polarization Features.
While the Stokes parameters $(\mathrm{S}_{0},\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3})$ provide a complete physical description of polarization, they are not always the most interpretable for human observers or visual analysis. To better convey polarization properties, several derived representations are commonly used. One of the most fundamental is the degree of polarization (DoP), which quantifies the fraction of total intensity carried by the polarized componenttyo2006review :
$\text{DoP}=\frac{\sqrt{S_{1}^{2}+S_{2}^{2}+S_{3}^{2}}}{S_{0}}=\sqrt{\left(\frac{\sqrt{S_{1}^{2}+S_{2}^{2}}}{S_{0}}\right)^{2}+\left(\frac{S_{3}}{S_{0}}\right)^{2}}=\sqrt{\text{DoLP}^{2}+\text{CoP}^{2}}$ (4)
This unified expression relates DoP to its two physically meaningful components: the degree of linear polarization (DoLP) and the circular polarization ratio (CoP), defined respectively as:
$\text{DoLP}=\frac{\sqrt{S_{1}^{2}+S_{2}^{2}}}{S_{0}},\quad\text{CoP}=\frac{S_{3}}{S_{0}}$ (5)
DoLP characterizes the proportion of linearly polarized light, while CoP indicates the contribution of circular polarization. In addition to magnitude, the angle of linear polarization (AoLP) describes the orientation of linear polarization and is defined as:
$\text{AoLP}=\frac{1}{2}\arctan\left(\frac{S_{2}}{S_{1}}\right)$ (6)
This formulation yields angles in the range $90^{\circ},90^{\circ}]$, representing the azimuthal angle of the polarization ellipse.
Together, these derived features—DoP, DoLP, CoP, and AoLP—offer more interpretable and visually meaningful descriptions of polarization than raw Stokes components. They also exhibit distinct gradient distributions: for example, AoLP often presents sharper local variations due to angular wrapping, whereas DoLP and CoP follow hyper-Laplacian-like distributionsjeon2024spectral . This highlights the need for feature-specific priors and visualization strategies when analyzing polarization images.
@@ 3.2 Benchmark Design
Next, we build our benchmark on the recent large-scale RGB-polarization dataset proposed by Jeon et al. jeon2024spectral . It provides high-quality, spatially aligned image pairs consisting of RGB inputs $\mathrm{S}_{0}$ and their corresponding Stokes components $\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3}]$, enabling supervised training. All images are resized to a fixed resolution of $H\times W$ before training. Both the RGB inputs and the Stokes targets are normalized to the range $[0,1]$ for consistent training across models.
@@
Following standard practices in low-level vision, we evaluate model performance using three complementary metrics: peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS). PSNR and SSIM assess pixel-level accuracy and structural fidelity, while LPIPS captures perceptual similarity in the feature space. All metrics are computed independently for each Stokes component and then averaged to report the overall performance. These metrics also align with the physical meaning of the Stokes components: high PSNR and SSIM and low LPIPS indicate that the predicted Stokes maps better preserve the spatial and structural polarization cues of the ground truth.
@@ 4 Experiments
In our experiments, we primarily adopt the dataset provided by Jeon et al. jeon2024spectral as the training and evaluation benchmark. Specifically, the first 1,000 RGB–Stokes image pairs are used for training, and the last 200 pairs from the dataset are reserved for testing. For a fair comparison, all image pairs are resized to $\times256$ for training and testing across all baseline models. In addition to the in-dataset evaluation, we further assess the generalization ability of the trained models on two external datasets: Qiu et al. qiu2021demos and Kurita et al. kurita2023sparse .
@@ 4.1 Implementation details
All experiments are conducted on a server equipped with four NVIDIA RTX A5000 GPUs, each with 24 GB of memory. To ensure reproducibility, we will release all code, model checkpoints, and configuration files upon acceptance. The following paragraphs describe the network architectures of all evaluated models, while detailed training hyperparameters are deferred to the Appendix.
Table 2: Quantitative results on RGB-to-polarization image estimation across datasets. We report average PSNR, SSIM, and LPIPS over Stokes components on Jeon et al. jeon2024spectral , Qiu et al. qiu2021demos , and Kurita et al. kurita2023sparse datasets. Higher PSNR/SSIM and lower LPIPS indicate better performance. We highlight the best and second-best results for each column.
Method Jeon et al. jeon2024spectral Qiu et al. qiu2021demos Kurita et al. kurita2023sparse PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS Wdiff ozdenizci2023restoring 13.11 0.6822 0.3772 11.44 0.6523 0.4042 11.56 0.6945 0.4311 DiT peebles2023scalable 23.33 0.8409 0.2447 14.74 0.7328 0.2829 17.86 0.8191 0.2874 RealFill tang2024realfill 21.81 0.8025 0.2654 15.19 0.7241 0.3028 18.09 0.8252 0.2654 Img2ImgTurbo parmar2024one 23.33 0.8725 0.3542 15.65 0.7566 0.3811 18.78 0.8504 0.3115 Restormer zamir2022restormer 23.98 0.8730 0.2369 14.77 0.5325 0.5186 18.85 0.8394 0.2697 Uformer wang2022uformer 24.34 0.8722 0.2267 14.68 0.7278 0.3160 18.74 0.8381 0.2627 MAE he2022masked 24.74 0.8876 0.2684 15.02 0.7401 0.3238 18.81 0.8499 0.2772
@@
Generation baselines
@@ 4.2 Quantitative evaluation
These quantitative trends can also be understood from the perspective of physical polarization properties. We find that estimating the $\mathrm{S}_{1}$ component is more difficult than $\mathrm{S}_{2}$ and $\mathrm{S}_{3}$. Physically, $\mathrm{S}_{1}$ represents the difference between horizontal and vertical polarization, which tends to be more sensitive to surface orientation and material properties. In many natural scenes, this component is weaker or more spatially uniform due to diffuse reflection, leading to a lower signal-to-noise ratio and making it harder for the model to learn accurate patterns from RGB inputs.
To further assess the generalization ability of RGB-to-polarization estimation, we evaluate models trained on the dataset provided by Jeon et al. jeon2024spectral using two additional benchmarks: Qiu et al. qiu2021demos and Kurita et al. kurita2023sparse . As shown in Table 2, the results exhibit consistent trends across datasets: restoration-based and pre-trained models (e.g., MAE, Uformer) generally achieve stronger performance, while diffusion-based models lag behind in quantitative metrics.
Figure 5: Qualitative comparison of predicted DoLP and AoLP maps, derived from Stokes components estimated from RGB input. Ground-truth maps are provided for reference, along with results from Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
@@ 4.3 Qualitative evaluation
Figures 3 and 4 show qualitative results of the estimated Stokes components generated by DiT peebles2023scalable , Img2ImgTurbo parmar2024one , Uformer wang2022uformer , and MAE he2022masked . Each column visualizes the reconstructed $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$ components produced by each method. The results reveal diverse reconstruction characteristics across models, including differences in texture sharpness, structural consistency, and polarization pattern styles. Overall, MAE and Uformer demonstrate superior visual quality compared to diffusion-based approaches. MAE is particularly effective in preserving fine structural details and material boundaries, with minimal artifacts across all Stokes components. However, the restored details from restoration-based methods still deviate from the correct results, indicating that such models may still lack the capacity to capture the underlying polarization cues precisely. In contrast, diffusion-based methods struggle to reconstruct the correct appearance and structure of polarization components, indicating difficulty in modeling these signals. Although Img2ImgTurbo achieves higher PSNR scores than DiT, its visual results are less faithful to the ground truth, particularly in terms of structural detail. This suggests a trade-off between pixel-level accuracy and perceptual quality among diffusion-based models.
Furthermore, Figures 5 and 6 visualize the corresponding DoLP and AoLP maps derived from the estimated Stokes components. These results provide an alternative perspective by directly reflecting the physical polarization cues. The restoration-based Uformer generally produces more accurate and stable DoLP patterns, whereas the diffusion-based DiT achieves relatively better AoLP maps, highlighting their complementary strengths in capturing polarization information.
@@
Figure 6: Qualitative comparison of predicted DoLP and AoLP maps, derived from Stokes components estimated from RGB input. Ground-truth maps are provided for reference, along with results from Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .